OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

OpenStudy (anonymous):

gradient = \[\frac{-1 - -7}{5 - 2} = \frac{6}{3} = 2\] then you do \[y - -7 = 2(x-5)\]

OpenStudy (anonymous):

Find slope first. Then, use y = mx + b equation to find b (the y-intercept).

OpenStudy (anonymous):

expanding and simplifying gives you y = 2x - 17